Use the elimination method to solve the system of equations.
a+9b=23
a+2b=2

Found 2 solutions by mananth, lynnlo:
Answer by mananth(16946)

1 a + 9 b = 23 .............1

1 a + 2 b = 2 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 9
-2 a -18 b = -46
9 a + 18 b = 18
Add the two equations
7 a = -28
/ 7
a = -4
plug value of a in (1)
1 a + 9 b = 23
-4 + 9 b = 23
9 b = 23 + 4
9 b = 27
b = 3
x= -4
y= 3
m.ananth@hotmail.ca

Answer by lynnlo(4176)